OpenStudy (anonymous):

What is an equation for the line that passes through points (–1, –4) and (1, 4)? Write the equation in slope-intercept form.

OpenStudy (anonymous):

y=4x

OpenStudy (anonymous):

the slope=(4-(-4))/(1-(-1))=(4+4)/(1+1)=8/2=4 y=4(x-1)+4 or y=4(x+1)-4 y=4x